OpenStudy (anonymous):

Write y = – 2/3 x + 7 in standard form using integers. 2x + 3y = 21 3x – 2y = 21 –2x – 3y = 21 –2x + 3y = 7

OpenStudy (phi):

standard form does not like a minus sign on the x value. That rules out two of your choices. (standard form puts the variables in alphabetical order, so first x then y) I would multiply your original equation by 3: every term on both sides of the = that should get you started.

OpenStudy (anonymous):

How do you multiple the -2/3x by 3 with the x there?

jimthompson5910 (jim_thompson5910):

3 times (-2/3x) = (3/1) times (-2/3x) 3 times (-2/3x) = (3*-2)/(1*3)x 3 times (-2/3x) = (-6)/(3)x 3 times (-2/3x) = -2x